Can you help me out with question 13 and 14. Thank you.

Here is the previous question it was referring to. N equals 10
Question 13 (4 points) Consider the research design described above. Does the study support her hypothesis? Compute the upper bound of the confidence interval using the following data: mean of the difference scores (subtract pretest from posttest): -1.5 standard error of the difference scores: 0.5 The formula for the Cl upper bound is [standard error of the difference scores]* [t critical value]+Imean of the difference scores Again, your answer must be expressed as a decimal out to 2 places, e.g., .20. Your Answer: Answer DView hint for Question 13 Question 14 (3 points) Would you reject the null hypothesis given your response to the question above?
Question 12 (1 point) Saved A researcher thinks that listening to classical music reduces anxiety. She measures the anxiety of 10 persons then plays Mozart's "Eine Kleine Nachtmusik" (listen here if you want--you know the tune!) for them. Following that the researcher measures their anxiety again. (Note that anxiety is measured on a scale from 1 to 7, with higher numbers indicating increased anxiety.) In this design, classical music is the independent variable criterion variable
Answer #1

Question 13:

n = Sample Size = 10

Mean of the difference scores = - 1.5

SE of difference scores = 0.5

Since Significance Level = \alpha not provided:
Take \alpha = 0.05

ndf= n - 1 = 10 - 1= 9

From Table critical values of t = \pm 2.2622

Upper bound of Confidence Interval:

- 1.5 + (2.2622 X 0.5)

= - 1.5 + 1.1311

= - 0.37

So,

Answer is:

- 0.37

Question 14:

Correct option:

Yes

Explanation:

Since all values in the Confidence Interval are negative, we conclude that the difference is significant. Reject null hypothesis.

Conclusion:
The data support the claim that listening to classical music reduces anxiety.